When Samuel was 7, Claire was 34. Claire’s age is now double Samuel’s age. How old will Samuel be in 10 years?
Answer by macston(5194)

.
S=Samuel's current age=C-27 yrs; C=Claire's current age=2S
.
C=2S
C=2(C-27 yrs)
C=2C-54 yrs
54 yrs=C
Claire is currently 54 years old.
.
S=C-27yrs=54-27 yrs=27 yrs
Samuel is currently 27 years old.
.
S+10 yrs=27 yrs + 10 yrs = 37 yrs
ANSWER: In ten years, Samuel will be 37 years old.